The hydrogen-filled airship Hindenburg (see Fig. 11-24) could cruise at 77 knots with the engines providing 4800 hp. Calculate the air drag force in newtons on the airship at this speed.

Grade:upto college level

1 Answers

Deepak Patra
8 years ago
To obtain air drag force in newtons, substitute 77 knots for v and 4800 hp for P in the equation F = P/v,
F = P/v
= 4800 hp/77 knots
= (4800 hp/77 knots) (745.7 W/1 hp) (1 knot/1.688 ft/s) (1 ft/s/0.3048 m/s)
= 9.0×104 W/m/s
= (9.0×104 W/m/s) (1 N/1 W/m/s)
=9.0×104 N …… (3)
From the above observation we conclude that, air drag force in newtons on the airship at this speed will be 9.0×104 N.